# A Connection Between Covariate Adjustment and Stratified Randomization in Randomized Clinical Trials

### v2

@@ 3.2 Connection with Covariate Adjustment
with asymptotic variance $\sigma^{2}(b+c)$ under simple randomization. According to Remark 1,
@@
As an example, consider the empirical estimator $\widehat{\delta}_{\text{emp}}=\widehat{\delta}_{\text{aug}}(0)$. With $b=0$, Remark 3 implies that $\sigma_{\text{st}}^{2}(0)=\min_{c\in\mathcal{C}}\sigma^{2}(c)$. Minimizing $\sigma^{2}(c)$ over $c\in\mathcal{C}$ is a special case of the efficient estimation problem considered in Section 2.2, with $\boldsymbol{W}$ replaced by $S$. Because $S$ is discrete, the minimum of $\sigma^{2}(c)$ is attained by the regression estimator based on model (1) with $\boldsymbol{V}=(I(S=2),\dots,I(S=K))^{\text{T}}$. With this choice of $\boldsymbol{V}$, model (1) is saturated and trivially correct, leading to a regression estimator that is efficient for $S$, which we denote by $\widehat{\delta}_{\text{reg}}^{S}$. Now, Remark 3 implies that the asymptotic variance of $\widehat{\delta}_{\text{emp}}$ under stratified randomization is equal to the asymptotic variance of $\widehat{\delta}_{\text{reg}}^{S}$ under simple randomization. This asymptotic equivalence makes intuitive sense. Stratified randomization produces balanced strata by design, while covariate adjustment can be regarded as an attempt to balance covariates by analysis. When covariate adjustment is focused on $S$ and optimally conducted, as in $\widehat{\delta}_{\text{reg}}^{S}$, it is expected to have the same asymptotic effect as stratified randomization based on $S$. It should be noted that this equivalence between covariate adjustment and stratified randomization is limited to the use of $S$. Covariate adjustment is able to incorporate information from additional baseline covariates of all types, whereas stratified randomization is constrained by the (typically small) number of strata that can be accommodated.
@@ 3.3 Practical Implications
In this subsection, we examine the implications of the results in Section 3.2 on clinical trial analysis and design. It is well recognized that stratified randomization should be acknowledged in the statistical analysis of trial data, which is usually interpreted to mean that treatment effect estimation should adjust for the stratification factor as a covariate 24, 25, 26. The recent FDA guidance on covariate adjustment 1 recommends that, under stratified randomization, “A covariate adjustment model should generally include strata variables and can also include covariates not used for stratifying randomization.” Accordingly, we will evaluate and compare the potential benefits of adjusting for the stratification factor and possibly other baseline covariates.
Let us start with the common practice of adjusting only for the stratification factor $S$. There are different ways to adjust for $S$, some of which may be more efficient than others. We will focus on the regression estimator $\widehat{\delta}_{\text{reg}}^{S}$, which makes optimal adjustment for $S$, as noted in Section 3.2. This implies that $\widehat{\delta}_{\text{reg}}^{S}$ is asymptotically equivalent to $\widehat{\delta}_{\text{aug}}(c_{\text{opt}})$, where $c_{\text{opt}}=\operatorname{arg\;min}_{c\in\mathcal{C}}\sigma^{2}(c)$ is the optimal augmentation function based on $S$ only. Under simple randomization, $\widehat{\delta}_{\text{reg}}^{S}$ is generally more efficient than $\widehat{\delta}_{\text{emp}}=\widehat{\delta}_{\text{aug}}(0)$ because $\sigma^{2}(c_{\text{opt}})=\min_{c\in\mathcal{C}}\sigma^{2}(c)\leq\sigma^{2}(0)$. Under stratified randomization, equation (6) implies that $\widehat{\delta}_{\text{emp}}$ has asymptotic variance $\sigma_{\text{st}}^{2}(0)=\min_{c\in\mathcal{C}}\sigma^{2}(0+c)=\sigma^{2}(c_{\text{opt}})$ and that $\widehat{\delta}_{\text{reg}}^{S}$ has asymptotic variance $\sigma_{\text{st}}^{2}(c_{\text{opt}})=\min_{c\in\mathcal{C}}\sigma^{2}(c_{\text{opt}}+c)=\sigma^{2}(c_{\text{opt}})$. Thus, under stratified randomization, $\widehat{\delta}_{\text{reg}}^{S}$ is as (in)efficient as $\widehat{\delta}_{\text{emp}}$ is. This is not surprising because the information in $S$ has already been used (optimally) in the randomization procedure. Only when we incorporate additional information from $\boldsymbol{W}$ into the analysis can we hope to improve efficiency over $\widehat{\delta}_{\text{emp}}$ under stratified randomization.
The information in $\boldsymbol{W}$ can be incorporated using a regression or augmented estimator described in Section 2.2. When all of $\boldsymbol{W}$ is included in covariate adjustment, the minimum attainable asymptotic variance is, according to Remark 2, $\sigma^{2}(b_{\text{opt}})=\min_{b\in\mathcal{B}}\sigma^{2}(b)$. Because $\mathcal{B}\supset\mathcal{C}$, this minimum is generally smaller than $\min_{c\in\mathcal{C}}\sigma^{2}(c)=\sigma^{2}(c_{\text{opt}})$, the minimum asymptotic variance that could be attained by adjusting for $S$ alone. The two minima are equal only when $\boldsymbol{W}$ contains no new prognostic information beyond $S$, which is unlikely given how $\boldsymbol{W}$ is chosen in the first place. Of course, $\sigma^{2}(b_{\text{opt}})$ is only a minimum and not necessarily the actual asymptotic variance of the estimator in use. However, in our numerical experience, it is difficult to find a situation in which an estimator that adjusts for all of $\boldsymbol{W}$ underperforms an estimator based on $S$, say $\widehat{\delta}_{\text{reg}}^{S}$.
@@
In designing a trial with stratified randomization, one often faces difficult choices in terms of which covariates to include in the stratification and how many strata to use. Including more covariates increases the potential benefit of stratified randomization, but may result in a large number of strata, some of which may be small and difficult to balance. Some biomarkers may be time-consuming to ascertain, and their inclusion in stratified randomization may cause a delay in randomization and treatment. Fortunately, Remark 3 indicates that the efficiency benefit of stratified randomization can be recovered through covariate adjustment. Therefore, it is desirable but not essential to include all important covariates in the stratification. To illustrate this point, we give some examples of covariate-adjusted estimators that have the same asymptotic variance under simple versus stratified randomization. The regression estimator $\widehat{\delta}_{\text{reg}}$ has this property if $\pi=1/2$ or if the vector $\boldsymbol{V}$ in model (1) includes the dummy variables $(I(S=2),\dots,I(S=K))$, according to Corollary 1 of Wang et al. 21. (There can, and should, be other covariates in $\boldsymbol{V}$ to improve efficiency.) A generic augmented estimator $\widehat{\delta}_{\text{aug}}(\widehat{b})$ can be modified to have this property using the following calibration technique proposed by Bannick et al. 22. For $k=1,\dots,K$, define
$\widehat{c}_{k}=\frac{\sum_{i=1}^{n}I(A_{i}=1,S_{i}=k)\widehat{\psi}_{(}\boldsymbol{O}_{i})}{\sum_{i=1}^{n}I(A_{i}=1,S_{i}=k)}-\frac{\sum_{i=1}^{n}I(A_{i}=0,S_{i}=k)\widehat{\psi}(\boldsymbol{O}_{i})}{\sum_{i=1}^{n}I(A_{i}=0,S_{i}=k)}-\frac{\sum_{i=1}^{n}I(S_{i}=k)\widehat{b}(\boldsymbol{W}_{i})}{\sum_{i=1}^{n}I(S_{i}=k)},$
@@
Remark 4.
@@
Under stratified randomization, the calibration technique does not improve efficiency (as can be argued along the lines of Remark 3) but can be useful for variance estimation. The variance reduction effect of stratified randomization implies that variance estimators developed for simple randomization may be biased upward under stratified randomization, unless the point estimator of $\delta$ has the same asymptotic variance under simple versus stratified randomization, as is the case for a calibrated estimator. For a generic augmented estimator $\widehat{\delta}_{\text{aug}}(\widehat{b})$ with $\widehat{b}\to b$, let $\widehat{\delta}_{\text{aug}}(\widehat{b}+\widehat{c})$ be defined as in the preceding paragraph; then $\sigma_{\text{st}}^{2}(b)$ is the common asymptotic variance for the two estimators under stratified randomization, and is consistently estimated by
@@
If $\widehat{b}$ is obtained using a machine learning method, the above variance estimator may suffer from a re-substitution bias in finite samples, which can be removed using cross-validation; see Section 2.3 of Zhang and Ma 10. For a non-augmented estimator, one may appeal to its asymptotic equivalence to an augmented estimator (say $\widehat{\delta}_{\text{aug}}(b)$), find a consistent estimator of $b$ (say $\widehat{b}$), and then apply the calibration technique to obtain a consistent variance estimator under stratified randomization.
@@ 4 Simulation
Let us first consider the case where $Y$ is continuous. In this case, the estimand $\delta$ is the mean difference between $Y(1)$ and $Y(0)$, and its true value is equal to 1 in each of the four scenarios described earlier. We compare six estimators of $\delta$: the empirical estimator $\widehat{\delta}_{\text{emp}}$, three regression estimators with different choices of $\boldsymbol{V}$, two augmented estimator (with or without calibration) where $\widehat{b}$ is obtained using a super learner 10. In the regression estimators, $h$ is the identity function (corresponding to ANCOVA) and $\boldsymbol{V}$ may be a vector of dummy variables for $S$, the original covariate vector $\boldsymbol{W}$, or $\boldsymbol{W}$ concatenated with dummy variables for $S$. The super learner in the augmented estimators is a linear combination (with coefficients optimized in five-fold cross-validation) of the following four algorithms: linear regression, linear regression with pairwise interactions, additive model, and recursive partitioning and regression tree. Under simple randomization, variance estimates are obtained using estimated influence functions, with the use of another five-fold cross-validation for the augmented estimators 10. Under stratified randomization, no adjustment in variance estimation is required for the regression estimators 21, and calibrated variance estimators (see Remark 4) are used for the empirical and augmented estimators. The different methods and designs (i.e., simple vs stratified randomization) are compared in terms of empirical bias, standard deviation, median standard error, and empirical coverage of 95% confidence intervals. The empirical method combined with simple randomization serves as a reference in efficiency comparisons. The relative efficiency of any method-design combination is defined as the inverse ratio of its empirical variance to that of the reference combination.
@@ 5 Example
We now illustrate our findings by analyzing real data from a National Institute of Neurological Disorders and Stroke (NINDS) trial of recombinant tissue plasminogen activator (rt-PA) for treating acute ischemic stroke 31. This trial enrolled 624 patients with acute ischemic stroke at eight sites, randomized them in a $1:1$ ratio to rt-PA or placebo within three hours of stroke onset, and evaluated clinical outcomes at 24 hours and three months post-treatment. This trial consisted of two (temporally ordered) parts, with 291 patients enrolled in part 1 and 333 in part 2. The main difference between the two parts was that the time from stroke onset to treatment (TOT) was within 90 minutes in part 1 and between 90 minutes and 180 minutes in part 2. Except for that difference, patients in the two parts were treated and evaluated in the same manner. Within each part, the randomization was stratified by site using permuted blocks of various sizes. In our retrospective analysis, we combine the two parts and consider the overall randomization scheme as stratified by site and study part.
@@
Following previous analyses of the NINDS trial data 31, 32, we compare treatments with respect to clinical outcomes at three months, as measured by the Barthel Index, the Modified Rankin Scale, the Glasgow Outcome Scale, and the National Institutes of Health Stroke Scale (NIHSS). The Barthel Index measures the ability to perform activities of daily living and yields a score that ranges from 0 to 100 (complete independence). The Modified Rankin Scale is a simplified overall assessment of function, with a score that ranges from 0 (absence of symptoms) to 5 (severe disability). The Glasgow Outcome Scale is a global assessment of function, with a score that ranges from 1 (good recovery) to 5 (death). The NIHSS is a 42-point scale with 0 indicating no stroke symptoms and higher scores for more severe symptoms. For ease of interpretation and comparison, each original score is linearly re-scaled to the unit interval in such a way that higher scores represent desirable outcomes. For each instrument, the efficacy of rt-PA relative to placebo is measured by the between-group difference (rt-PA minus placebo) in the re-scaled mean score.
@@
In our analysis, the covariate vector $\boldsymbol{W}$ consists of continuous TOT, baseline NIHSS, age, preexisting disability, and history of diabetes. (Study part is not included in $\boldsymbol{W}$ because it is determined by TOT, and clinical site is not included because some sites have few patients.) Six patients with missing covariate data are excluded from this analysis. The remaining data is analyzed using the empirical method, the regression method based on $\boldsymbol{W}$, and the augmentation method based on $\boldsymbol{W}$ with or without calibration. (It is infeasible to adjust for strata in the regression method due to data sparsity.) The methods are identical to those compared in the simulation study (for a continuous outcome), except that the super learner library includes random forest in addition to the algorithms described in Section 4.

### v3

@@ 3.2.2 Analytical analogue of stratified randomization
with asymptotic variance $\sigma^{2}(b+c)$ under simple randomization. According to Proposition 1,
@@
3.2.3 Asymptotic equivalence between stratified randomization and analysis
Consider the empirical estimator $\widehat{\delta}_{\text{emp}}=\widehat{\delta}_{\text{aug}}(0)$. With $b=0$, equation (4) implies that
$\sigma_{\text{st}}^{2}(0)=\min_{c\in\mathcal{C}}\sigma^{2}(c).$ (5)
Minimizing $\sigma^{2}(c)$ over $c\in\mathcal{C}$ is a special case of the efficient estimation problem considered in Section 2.2, with $\boldsymbol{W}$ replaced by $S$. Because $S$ is discrete, the minimum of $\sigma^{2}(c)$ is attained by the regression estimator (described in Appendix A) based on model (A.1) with $\boldsymbol{V}=(I(S=2),\dots,I(S=K))^{\text{T}}$. With this choice of $\boldsymbol{V}$, model (A.1) is saturated and trivially correct, leading to a regression estimator that is efficient for $S$, which we denote by $\widehat{\delta}_{\text{reg}}^{S}$. Now, equation (5) indicates that the asymptotic variance of $\widehat{\delta}_{\text{emp}}$ under stratified randomization is equal to the asymptotic variance of $\widehat{\delta}_{\text{reg}}^{S}$ under simple randomization. This result coincides with Corollary 2 of Ye et al. 20 when model (A.1) has an identity link. This asymptotic equivalence makes intuitive sense. Stratified randomization produces balanced strata by design, while a stratified analysis attempts to balance covariates by analysis. When the latter is optimally conducted as in $\widehat{\delta}_{\text{reg}}^{S}$, it is expected to have the same asymptotic effect as stratified randomization based on $S$.
It should be noted that this asymptotic equivalence between design and analysis is limited to the use of strata information. The stratified analysis based on $\widehat{\delta}_{\text{reg}}^{S}$ is a special case of covariate adjustment, which is more general and able to incorporate information from additional baseline covariates in the analysis stage. In contrast, stratified randomization is constrained by the (typically small) number of strata that can be accommodated.
@@ 3.3 Practical Implications
3.3.1 Trial design
@@
In designing a trial with stratified randomization, one often faces difficult choices in terms of which covariates to include in the stratification and how many strata to use. Including more covariates increases the potential benefit of stratified randomization, but may result in a large number of strata, some of which may be small and difficult to balance. Some biomarkers may be time-consuming to ascertain, and their inclusion in stratified randomization may cause a delay in randomization and treatment. Fortunately, equation (4) indicates that the efficiency benefit of stratified randomization can be recovered through covariate adjustment. Therefore, it is desirable but not essential to include all important covariates in the stratification.
To illustrate this point, we give some examples of covariate-adjusted estimators that have the same asymptotic variance under simple versus stratified randomization. The regression estimator $\widehat{\delta}_{\text{reg}}$ (described in Appendix A) has this property if the vector $\boldsymbol{V}$ in model (A.1) includes the dummy variables $(I(S=2),\dots,I(S=K))$, according to Theorem 2 of Ye et al. 20 and Corollary 1 of Wang et al. 21. (There can, and should, be other covariates in $\boldsymbol{V}$ to improve efficiency.) A generic augmented estimator $\widehat{\delta}_{\text{aug}}(\widehat{b})$ can be modified to have this property using the following calibration technique proposed by Bannick et al. 22. For $k=1,\dots,K$, define
$\widehat{c}_{k}=\frac{\sum_{i=1}^{n}I(A_{i}=1,S_{i}=k)\widehat{\psi}(\boldsymbol{O}_{i})}{\sum_{i=1}^{n}I(A_{i}=1,S_{i}=k)}-\frac{\sum_{i=1}^{n}I(A_{i}=0,S_{i}=k)\widehat{\psi}(\boldsymbol{O}_{i})}{\sum_{i=1}^{n}I(A_{i}=0,S_{i}=k)}-\frac{\sum_{i=1}^{n}I(S_{i}=k)\widehat{b}(\boldsymbol{W}_{i})}{\sum_{i=1}^{n}I(S_{i}=k)},$
@@
3.3.2 Data analysis
It is well recognized that stratified randomization should be acknowledged in the statistical analysis of trial data, which is usually interpreted to mean that treatment effect estimation should adjust for the stratification factor as a covariate 28, 29, 30. The recent FDA guidance on covariate adjustment 1 recommends that, under stratified randomization, “A covariate adjustment model should generally include strata variables and can also include covariates not used for stratifying randomization.” Accordingly, we will evaluate and compare the potential benefits of adjusting for the stratification factor and possibly other baseline covariates under stratified randomization.
Let us start with the common practice of adjusting only for the stratification factor $S$. There are different ways to adjust for $S$, some of which may be more efficient than others. We will focus on the regression estimator $\widehat{\delta}_{\text{reg}}^{S}$, which makes optimal adjustment for $S$, as noted in Section 3.2.3. This implies that $\widehat{\delta}_{\text{reg}}^{S}$ is asymptotically equivalent to $\widehat{\delta}_{\text{aug}}(c_{\text{opt}})$, where $c_{\text{opt}}=\operatorname{arg\;min}_{c\in\mathcal{C}}\sigma^{2}(c)$ is the optimal augmentation function based on $S$ only (see Figure 1 for an illustration). Under simple randomization, $\widehat{\delta}_{\text{reg}}^{S}$ is generally more efficient than $\widehat{\delta}_{\text{emp}}=\widehat{\delta}_{\text{aug}}(0)$ because $\sigma^{2}(c_{\text{opt}})=\min_{c\in\mathcal{C}}\sigma^{2}(c)\leq\sigma^{2}(0)$. Under stratified randomization, equation (4) implies that $\widehat{\delta}_{\text{emp}}$ has asymptotic variance $\sigma_{\text{st}}^{2}(0)=\min_{c\in\mathcal{C}}\sigma^{2}(0+c)=\sigma^{2}(c_{\text{opt}})$ and that $\widehat{\delta}_{\text{reg}}^{S}$ has asymptotic variance $\sigma_{\text{st}}^{2}(c_{\text{opt}})=\min_{c\in\mathcal{C}}\sigma^{2}(c_{\text{opt}}+c)=\sigma^{2}(c_{\text{opt}})$. Thus, under stratified randomization, $\widehat{\delta}_{\text{reg}}^{S}$ is as (in)efficient as $\widehat{\delta}_{\text{emp}}$ is. This is not surprising because the information in $S$ has already been used (optimally, according to equation (4)) in the randomization procedure. Only when we incorporate additional information from $\boldsymbol{W}$ into the analysis can we hope to improve efficiency over $\widehat{\delta}_{\text{emp}}$ under stratified randomization.
The information in $\boldsymbol{W}$ can be incorporated using a regression estimator or an augmented estimator described in Section 2.2. When all of $\boldsymbol{W}$ is included in covariate adjustment, the minimum attainable asymptotic variance is $\sigma^{2}(b_{\text{opt}})=\min_{b\in\mathcal{B}}\sigma^{2}(b)$, as noted in Section 3.2.1. Because $\mathcal{B}\supset\mathcal{C}$, this minimum is generally smaller than $\min_{c\in\mathcal{C}}\sigma^{2}(c)=\sigma^{2}(c_{\text{opt}})$, the minimum asymptotic variance that could be attained by adjusting for $S$ alone. The two minima are equal only when $\boldsymbol{W}$ contains no new prognostic information beyond $S$, which is unlikely given how $\boldsymbol{W}$ is chosen in the first place. Of course, $\sigma^{2}(b_{\text{opt}})$ is only a minimum and not necessarily the actual asymptotic variance of the estimator in use. However, in our numerical experience, it has been difficult to find a situation in which an estimator that adjusts for all of $\boldsymbol{W}$ underperforms an estimator based on $S$, say $\widehat{\delta}_{\text{reg}}^{S}$.
As noted in Section 3.2.1, under both simple and stratified randomization, the key to efficient estimation is to adjust for all of $\boldsymbol{W}$ using an appropriate model or method with a limiting augmentation function close to $b_{\text{opt}}$. From this perspective, it is not clear why, under stratified randomization, the stratification factor should take precedence over other baseline covariates in covariate adjustment as suggested by regulatory guidelines 28, 29, 30, 1. This is not to argue against adjusting for $S$ but rather to question the rationale for prioritizing $S$ over other covariates. It is possible that, for a particular model or method, adding dummy variables for $S$ to the covariate vector $\boldsymbol{W}$ may actually improve estimation precision. If this happens, it is because the model or method for covariate adjustment does not make effective use of $\boldsymbol{W}$ (to which $S$ adds no new information), not because $S$ is the stratification factor. If anything, the use of $S$ in stratified randomization should lead to a de-prioritization of $S$ in covariate adjustment, because stratified randomization already makes use of some of the information in $S$, as shown in equation (4).
@@ 3.3.2 Data analysis
The variance reduction effect of stratified randomization implies that variance estimators developed for simple randomization may be biased upward under stratified randomization, unless the point estimator of $\delta$ has the same asymptotic variance under simple versus stratified randomization. Section 3.3.1 gives examples of estimators that do not receive a variance reduction from stratified randomization, including a calibrated estimator that turns out useful for variance estimation. For a generic augmented estimator $\widehat{\delta}_{\text{aug}}(\widehat{b})$ with $\widehat{b}\to b$, let $\widehat{\delta}_{\text{aug}}(\widehat{b}+\widehat{c})$ be the calibrated estimator defined in Section 3.3.1; then $\sigma_{\text{st}}^{2}(b)$ is the common asymptotic variance of both $\widehat{\delta}_{\text{aug}}(\widehat{b})$ and $\widehat{\delta}_{\text{aug}}(\widehat{b}+\widehat{c})$ under stratified randomization, and is consistently estimated by
@@
If $\widehat{b}$ is obtained using a machine learning method, the above variance estimator may suffer from a re-substitution bias in finite samples, which can be removed using cross-validation (see Section 2.3 of Zhang and Ma 10). For a non-augmented estimator, one may appeal to its asymptotic equivalence to an augmented estimator (say $\widehat{\delta}_{\text{aug}}(b)$), find a consistent estimator of $b$ (say $\widehat{b}$), and then apply the calibration technique to obtain a consistent variance estimator under stratified randomization.
Interestingly, variance estimation provides one possible reason for including $S$ in covariate adjustment. As noted in Section 3.3.1, the regression estimator $\widehat{\delta}_{\text{reg}}$ does not require a bias correction in variance estimation due to stratified randomization if the covariate vector $\boldsymbol{V}$ in the model includes dummy variables for $S$. It is thus convenient for variance estimation to utilize such a regression estimator with such a covariate configuration. This convenience may be of practical importance but seems insufficient to support the recommendation of always adjusting for $S$ as a profound general principle, because the convenience argument has some limitations. First, it is limited to a specific type of regression estimators and not applicable to general (and potentially more efficient) augmented estimators. Second, as noted in the preceding paragraph, it is straightforward to obtain a valid variance estimator for a generic treatment effect estimator, and computational software is also publicly available 34. In light of these limitations, it remains unclear what the rationale is for the general recommendation of always adjusting for $S$ in the analysis following stratified randomization.
@@ 4 Simulation
Let us first consider the case where $Y$ is continuous. In this case, the estimand $\delta$ is the mean difference between $Y(1)$ and $Y(0)$, and its true value is equal to 1 in each of the four scenarios described earlier. We compare six estimators of $\delta$: the empirical estimator $\widehat{\delta}_{\text{emp}}$, three regression estimators with different choices of $\boldsymbol{V}$, two augmented estimator (with or without calibration) where $\widehat{b}$ is obtained using a super learner 10. In the regression estimators, $h$ is the identity function (corresponding to ANCOVA) and $\boldsymbol{V}$ may be a vector of dummy variables for $S$, the original covariate vector $\boldsymbol{W}$, or $\boldsymbol{W}$ concatenated with dummy variables for $S$. The super learner in the augmented estimators is a linear combination (with coefficients optimized in five-fold cross-validation) of the following four algorithms: linear regression, linear regression with pairwise interactions, additive model, and recursive partitioning and regression tree. Under simple randomization, variance estimates are obtained using estimated influence functions, with the use of another five-fold cross-validation for the augmented estimators 10. Under stratified randomization, no adjustment in variance estimation is required for the regression estimators that explicitly adjust for $S$, and calibrated variance estimators (see Section 3.3.2) are used for the other estimators. The different methods and designs (i.e., simple vs stratified randomization) are compared in terms of empirical bias, standard deviation, median standard error, and empirical coverage of 95% confidence intervals. The empirical method combined with simple randomization serves as a reference in efficiency comparisons. The relative efficiency of any method-design combination is defined as the inverse ratio of its empirical variance to that of the reference combination.
@@ 5 Example
We now illustrate our findings by analyzing real data from a National Institute of Neurological Disorders and Stroke (NINDS) trial of recombinant tissue plasminogen activator (rt-PA) for treating acute ischemic stroke 35. This trial enrolled 624 patients with acute ischemic stroke at eight sites, randomized them in a $1:1$ ratio to rt-PA or placebo within three hours of stroke onset, and evaluated clinical outcomes at 24 hours and three months post-treatment. This trial consisted of two (temporally ordered) parts, with 291 patients enrolled in part 1 and 333 in part 2. The main difference between the two parts was that the time from stroke onset to treatment (TOT) was within 90 minutes in part 1 and between 90 minutes and 180 minutes in part 2. Except for that difference, patients in the two parts were treated and evaluated in the same manner. Within each part, the randomization was stratified by site using permuted blocks of various sizes. In our retrospective analysis, we combine the two parts and consider the overall randomization scheme as stratified by site and study part. Our analysis excludes one site which enrolled only one patient (while the other sites enrolled at least 39 patients per site).
@@
Following previous analyses of the NINDS trial data 35, 36, we compare treatments with respect to clinical outcomes at three months, as measured by the Barthel Index, the Modified Rankin Scale, the Glasgow Outcome Scale, and the National Institutes of Health Stroke Scale (NIHSS). The Barthel Index measures the ability to perform activities of daily living and yields a score that ranges from 0 to 100 (complete independence). The Modified Rankin Scale is a simplified overall assessment of function, with a score that ranges from 0 (absence of symptoms) to 5 (severe disability). The Glasgow Outcome Scale is a global assessment of function, with a score that ranges from 1 (good recovery) to 5 (death). The NIHSS is a 42-point scale with 0 indicating no stroke symptoms and higher scores for more severe symptoms. For ease of interpretation and comparison, each original score is linearly re-scaled to the unit interval in such a way that higher scores represent desirable outcomes. For each instrument, the efficacy of rt-PA relative to placebo is measured by the between-group difference (rt-PA minus placebo) in the re-scaled mean score.
@@
In our analysis, the covariate vector $\boldsymbol{W}$ consists of dummy variables for site, continuous TOT, baseline NIHSS, age, preexisting disability, and history of diabetes. (Study part is not included in $\boldsymbol{W}$ because it is determined by TOT.) Six patients with missing covariate data are excluded from this analysis. The remaining data is analyzed using the empirical method, three regression methods based on $S$ and/or $\boldsymbol{W}$, and the augmentation method based on $\boldsymbol{W}$ with or without calibration. The methods are identical to those compared in the simulation study (for a continuous outcome), except that the super learner library includes random forest in addition to the algorithms described in Section 4.